In each problem, make sure that you are clearly defining random variables, stating their distributions, and writing down the formulas that you are using. (That is, write down the pmf, write down mean and variance formulas.)

The number of fatalities due to shark attacks during a year is modeled using a Poisson distribution. (Rigorous justification: Sharks are fish.) The International Shark Attack File (ISAF) reports that during a 5 year period, there are on average 4.

4 fatalities per year due to shark attacks.

Assume that this average will remain the same in the future, and find the following probabilities regarding shark attacks in the next 5 years:

a. No shark fatalities recorded in a given single year.

b. Sharks cause at least six human deaths in a given single year.

c. No shark fatalities will occur during the next 5 year period.

d. At most 12 shark fatalities will occur during the next 5 year period.
Answer #1

Ans:

Poisson random variable with mean rate=4 per year

a)

P(x=0)=e-4=0.0183

b)

P(x>=6)=1-P(x<=5)=1-Poisson(5,4,true)=0.2149

c)Average number of fatalities in 5 years=4*5=20

P(x=0)=e-20=0.0000

d)

P(x<=12)=Poisson(12,20,true)=0.0390